OpenStudy (anonymous):

sin(arctan(4/3)) I'll draw the triangle below. part 1: Find the measure of the hypotenuse of this triangle. part 2: Angle theta is the angle whose tangent is 4/3. Find the sine of angle theta.

OpenStudy (anonymous):



OpenStudy (psymon):

Right, it wants to know the sine of a triangle that has a tangent of 4/3. So now you just need to find the hypotenuse by pythagorean theorem.

OpenStudy (anonymous):

oh my gosh, how did I not know that!? Wow that was such an obvious answer! haha! I was totally thinking of some complicated method!

OpenStudy (psymon):

Lol, yeah, nothing crazy xD

OpenStudy (anonymous):

I know what to do now for the next part, thanks so much! :)

OpenStudy (ybarrap):

This is a 3-4-5 triangle. Now that you know hypotenuse, Solve:\( \tt 5\sin)\theta)=4 \) for \( sin(\theta\)) and you're done.